OpenStudy (anonymous):

Can someone help me?!?!!? What is the product of 8 10 --- (M) --- (m is multiplying) 15 11 in simplest form? 16 ---- 33 ------------------------ 9 --- 13 ------------------------ 88 --- 150 ------------------------ 16 ---- 165

OpenStudy (mathteacher1729):

Do you mean \[\huge \frac{8}{10}\cdot \frac{15}{11}\]

OpenStudy (anonymous):

no :/ its 8/15 then 10/11

OpenStudy (mathteacher1729):

\[\huge \frac{8}{15}\cdot \frac{10}{11}\]

OpenStudy (anonymous):

Yea!! Thats it^^

OpenStudy (anonymous):